# GGTalker: Talking Head Systhesis with Generalizable Gaussian Priors and Identity-Specific Adaptation

### v1

@@ 1 Introduction
With the increasing demand for virtual reality [24], digital agents [33] and other applications, creating realistic, vivid and expressive talking heads has become a crucial task of computer vision. Significant progress has been made in this area, with numerous remarkable achievements. In particular, with the development of efficient 3D rendering technologies such as Neural Radiance Fields (NeRF) [23] and 3D Gaussian Splatting (3DGS) [14] in recent years, 3D-based talking heads [11, 18, 25, 38, 19, 7, 1, 39] are gradually gaining more attention than 2D-based ones [26, 43, 30, 6, 41, 29], thanks to their strong identity consistency and fast rendering speed.
@@
Specifically, to introduce facial shape priors and explicitly control head poses, we use FLAME [21] as the intermediate representation for the talking head and rig the Gaussians to the FLAME mesh [28]. To learn Audio-Expression Priors, we train a diffusion transformer model on large-scale datasets to predict audio-conditioned expressions. For Expression-Visual Priors, we leverage the UV layout of the canonical mesh to elegantly predict 3D priors in the UV space. We use an Identity-Gaussian Generator to predict reasonable initial 3D Gaussian head from a single reference image and introduce a source-target self-supervised training paradigm to accommodate cross-identity monocular and multi-view datasets. To adapt to the target individual’s speaking style and texture details, we jointly fine-tune the lip motion prior and texture prior on specific data. Simultaneously, we optimize the FLAME parameters to mitigate the errors caused by the current imperfect monocular tracking algorithms. We further introduce a color MLP to generate motion-aligned sharp textures and a Body Inpainter to blend rendered results with the torso and background, producing indistinguishable, photorealistic video results. Comprehensive experiments show that GGTalker outperforms in rendering quality, 3D consistency, lip-sync accuracy, training efficiency and inference speed.
@@ 2 Related Work
2D-based methods, including Generative Adversarial Networks (GAN) and diffusion models, are commonly used for synthesizing talking heads due to their powerful generative capabilities [46, 31, 22]. For instance, Wav2Lip [26] introduces a lip synchronization expert to supervise lip movements, enforcing the consistency of lip movements with the audio. IP-LAP [45] proposes a two-stage framework consisting of audio-to-landmark generation and landmark-to-video rendering procedures. However, these methods generate only the mouth region or the low half of face, which can lead to uncoordinated facial movements and artifacts. Apart from video-based methods, some approaches enable speech-driven animation from a single image [41, 36, 5]. For example, Hallo [36] introduces a hierarchical cross-attention mechanism to augment the correlation between audio inputs and non-identity-related motions. However, relying solely on a single reference image making it difficult to maintain the identity consistently and talking style.

### v2

@@ 1 Introduction
With the increasing demand for virtual reality [25], digital agents [37] and other applications, creating realistic, vivid and expressive talking heads has become a crucial task of computer vision. Significant progress has been made in this area, with numerous remarkable achievements. In particular, with the development of efficient 3D rendering technologies such as Neural Radiance Fields (NeRF) [24] and 3D Gaussian Splatting (3DGS) [14] in recent years, 3D-based talking heads [11, 18, 26, 42, 19, 7, 1, 45, 28, 21, 20, 44, 43] are gradually gaining more attention than 2D-based ones [30, 49, 34, 6, 47, 33], thanks to their strong identity consistency and fast rendering speed.
@@
Specifically, to introduce facial shape priors and explicitly control head poses, we use FLAME [22] as the intermediate representation for the talking head [27] and rig the Gaussians to the FLAME mesh [32]. To learn Audio-Expression Priors, we train a diffusion transformer model on large-scale datasets to predict audio-conditioned expressions. For Expression-Visual Priors, we leverage the UV layout of the canonical mesh to elegantly predict 3D priors in the UV space. We use an Identity-Gaussian Generator to predict reasonable initial 3D Gaussian head from a single reference image and introduce a source-target self-supervised training paradigm to accommodate cross-identity monocular and multi-view datasets. To adapt to the target individual’s speaking style and texture details, we jointly fine-tune the lip motion prior and texture prior on specific data. Simultaneously, we optimize the FLAME parameters to mitigate the errors caused by the current imperfect monocular tracking algorithms. We further introduce a color MLP to generate motion-aligned sharp textures and a Body Inpainter to blend rendered results with the torso and background, producing indistinguishable, photorealistic video results. Comprehensive experiments show that GGTalker outperforms in rendering quality, 3D consistency, lip-sync accuracy, training efficiency and inference speed.
@@ 2 Related Work
2D-based methods, including Generative Adversarial Networks (GAN) and diffusion models, are commonly used for synthesizing talking heads due to their powerful generative capabilities [52, 35, 23, 29]. For instance, Wav2Lip [30] introduces a lip synchronization expert to supervise lip movements, enforcing the consistency of lip movements with the audio. IP-LAP [51] proposes a two-stage framework consisting of audio-to-landmark generation and landmark-to-video rendering procedures. However, these methods generate only the mouth region or the low half of face, which can lead to uncoordinated facial movements and artifacts. Apart from video-based methods, some approaches enable speech-driven animation from a single image [47, 40, 5]. For example, Hallo [40] introduces a hierarchical cross-attention mechanism to augment the correlation between audio inputs and non-identity-related motions. However, relying solely on a single reference image making it difficult to maintain the identity consistently and talking style.